# Probabilistic Kernel Function for Fast Angle Testing

### v2

@@ 4 Analysis of Probability Guarantees
(2) Let $\bm{q}$, $\bm{v_{1}}$ and $\bm{v_{2}}$ be three normalized vectors on $\mathbb{S}^{d-1}$ such that $\left\langle\bm{q},\bm{v_{1}}\right\rangle>\left\langle\bm{q},\bm{v_{2}}\right\rangle$. The probability $\mathbb{P}[K^{1}_{S}(\bm{q},\bm{v_{1}})>K^{1}_{S}(\bm{q},\bm{v_{2}})|A_{S}(\bm{q})=\cos\psi]$ increases as $\psi$ decreases in $(0,\pi)$. In particular, when $\psi\in(0,\pi/2)$, $P[K^{1}_{S}(\bm{q},\bm{v_{1}})>K^{1}_{S}(\bm{q},\bm{v_{2}})|A_{S}(\bm{q})=\cos\psi]>0.5$, that is, $K^{1}_{S}$ satisfies the probability guarantee in Problem 1.1.
@@ 5 Implementation and Complexity Analysis
We discuss how to configure $S$, and then analyze the complexity of $K^{1}_{S}$ and $K^{2}_{S}$. Based on the discussion in Sec. 4, we observe that small reference angles are preferred. Thus, given $m$, our goal is to construct a set $S$ of $m$ points on $\mathbb{S}^{d-1}$, denoted by $S_{m}$, such that the reference angle $A_{S_{m}}(\cdot)$ is minimized. Due to the effect of the random rotation matrix $H$, the optimal configurations denoted by $\bar{S}_{m}$ and ${S}_{m}^{*}$, can be obtained either in the sense of expectation or in the sense of the worst case, respectively:
@@
where $U(\mathbb{S}^{d-1})$ denotes the uniform distribution on the sphere. By the definitions of $\bar{S}_{m}$ and ${S}_{m}^{\ast}$, they correspond to the configurations that achieve the smallest expected value and the smallest maximum value of $A_{S}(\bm{v})$, respectively. On the other hand, finding the exact solutions for $\bar{S}_{m}$ and ${S}_{m}^{\ast}$ is closely related to the best covering problem, which is highly challenging and remains open in the general case. To the best of the authors’ knowledge, the optimal configuration ${S}_{m}^{\ast}$ is only known when $m\leq d+3$ (Borodachov et al., 2019). In light of this, we provide two configurations of $S$: one relies on random antipodal projections (Alg. 1), and the other is built using multiple cross-polytopes (Alg. 2). Each has its own advantages. Alg. 1 enables the estimation of reference angles, while Alg. 2 can empirically produce slightly smaller reference angles and is more efficient for projection computation.
@@
2   Generate $m/2$ points along with their antipodal points i.i.d. on $S^{d^{\prime}-1}$  Scale the norm of all $m$ points in this iteration to $1/\sqrt{L}$, and collect the vectors after scaling
@@
Generate $N$ points randomly and independently on $\mathbb{S}^{d-1}$, where $N$ is a sufficiently large number
@@ 5      Repeat the above iteration and select $b/2$ antipodal pairs from the rotated cross-polytope
6  For the generated $S\in\mathbb{S}^{d^{\prime}-1}$, compute $\tilde{J}(S,N)$ and maintain the largest $S$ denoted by $S_{\text{max}}$
@@
Before proceeding into the detail of algorithms, we introduce a quantity $J(S)$ as follows.
$J(S)=\mathbb{E}_{\bm{v}\in U(\mathbb{S}^{d-1})}[(A_{S}(\bm{v}))].$ (7)
By definition, $J(S)$ denotes the expected value of the cosine of the reference angles w.r.t $S$. This quantity is consistent with our theory, as a random rotation is applied to $\bm{v}$ or $\bm{q}$ in Eq. (2) and Eq. (3). Based on the previous discussion, for a fixed $m$, $J(S)$ is minimized when $S=\bar{S}_{m}$, which is hard to compute. Thus, we take $N$ to be sufficiently large and let $v_{1,N},\ldots,v_{N,N}$ be vectors drawn independently and uniformly from $U(\mathbb{S}^{d-1})$. Let $\tilde{J}(S,N)=[\Sigma_{i=1}^{N}A_{S}(\bm{v_{i,N}})]/N$. By the law of large number, we can approximate $J(S)$ by $\tilde{J}(S,N)$ when $N$ is sufficiently large.
@@
(1) (Utilization of antipodal pairs and cross-polytopes) We use the antipodal pair or the cross-polytope as our building block for the following three reasons. (i) Since all the projection vectors are antipodal pairs, the evaluation time of projection vectors can be halved. (ii) Both of two structures can ensure that the assumption $A_{S}(\bm{v})>0$ holds, such that the condition in Lemma 4.3 is always satisfied. (iii) The result in Borodachov (2022) shows that, for $m=2d$, under mild conditions, the $2d$ vertices of a cross-polytope can be proven to have the smallest covering radius, that is, the smallest reference angle in the worst case. Although the results in the case $m>2d$ are unknown, we can rotate the fixed cross-polytope in random directions to generate multiple cross-polytopes until we obtain $m$ vectors, which explains the steps from 3 to 7 in Alg. 2.
@@
(3) (Accuracy boosting via multiple levels) Increasing $m$ can lead to a smaller reference angle. The analysis in Lu et al. (2024) shows that, for certain angle-thresholding problems requiring high accuracy, an exponential increase in $m$, rather than a linear one, can be effective. Therefore, similar to Lu et al. (2024), we use a product-quantization-like technique (Jégou et al., 2011) to partition the original space into $L$ subspaces (levels), which is adopted in both algorithms. By concatenating equal-length sub-vectors from these $L$ subspaces, we can virtually generate $m^{L}$ normalized projected vectors. As will be shown in Lemma B.1, the introduction of $L$ can significantly decrease the reference angle.
(4) (Fast projection computation via multi-cross-polytypes) In Eq. (2) and Eq. (3), we need to compute $HS\bm{v}$ in the indexing phase. If the projection time is a concern in practice, we can use the Fast Johnson–Lindenstrauss transformation to accelerate this process. Specifically, we use Alg. 2 with $R=1$. When $L=1$, the cost of computing $HS\bm{v}$ is $O(\max(d,m)\log d)$. When $L>1$, the cost of computing $HS\bm{v}$ can be reduced to $O(d\log d+mL\log(d/L))$.
@@
In the indexing phase, for a fixed dataset $\mathcal{D}$ of size $n$, the complexities of $K^{1}_{S}$ and $K^{2}_{S}$ are $O(nmd)$ and $O(nd\log d+nmd)$, respectively. In the query phase, $K^{1}_{S}$ requires $O(md)$- time for random projection, and for each $v\in\mathcal{D}$, it spends $O(1)$-time for computing $K^{1}_{S}(q,v)$, while $K^{2}_{S}$ spends $O(d\log d+md)$-time for random projection and random rotation, and for each $v\in D$, it spends $O(L)$-time for computing $K^{2}_{S}(q,v)$, where $L\ll d$. Particularly, if $S=S_{\text{pol}}(m,L)$ with $R=1$, the complexities of $K^{1}_{S}$ and $K^{2}_{S}$ in the indexing phase can be reduced to $O(\max(d,m)n\log d)$ and $O(nd\log d+nmL\log(d/L))$, respectively, where $m$ is greater than $d/L$ by user-specification.
@@ 6.1 Improvement on CEOs-Based Techniques
As for $K^{1}_{S}$, we can use it to improve CEOs, which is used for MIPS and further applied to accelerate LSH-based ANNS (Andoni et al., 2015) and DBSCAN (Xu & Pham, 2024). Since CEOs is originally designed for inner products, we generalize $K^{1}_{S}$ to $K^{1^{\prime}}_{S}$ as follows to align with CEOs:
@@
It is easy to see that, with two minor modifications, that is, replacing $\bm{v}\in\mathbb{S}^{d-1}$ with $\bm{v}\in\mathbb{R}^{d}$, and replacing $x$ with $\|\bm{v}\|x$ in Eq. (4), Lemma 4.2 still holds. Therefore, $K^{1^{\prime}}_{S}$ can be regarded as a reasonable kernel function for inner products. Then, we can apply $K^{1^{\prime}}_{S}$ to the algorithm in Pham (2021); Pham & Liu (2022); Xu & Pham (2024). We only need to make the following modification. In these algorithms, the random Gaussian matrix, which denotes the set of projection vectors, can be replaced by $S_{\text{sym}}$ or $S_{\text{pol}}$, with the other parts unchanged. This substitution does not change the complexity of the original algorithms. To distinguish this projection technique based on $K^{1^{\prime}}_{S}$ from CEOs, we refer to it as KS1 (see Alg. 4 for the projection structure of KS1). In the experiments, we will demonstrate that KS1 yields a slight improvement in recall rates over CEOs, owing to a smaller reference angle.
@@ 6.2 A New Probabilistic Test in Similarity Graph
Lu et al. (2024) proposed a $(\delta,1-\epsilon)$-routing test called PEOs test. Based on $K^{2}_{S}$, we propose a new routing test for $\ell_{2}$ distance, called the KS2 test, as follows (see Sec. C.1 in Appendix for more detail).
$\Sigma^{L}_{i=1}{\bm{{q_{i}}^{\top}u^{i}_{e[i]}}}\geq A_{S}(\bm{v})\cdot\frac{\|\bm{w}\|^{2}/2-\tau-\bm{v}^{\top}\bm{q}}{\|\bm{e}\|}.$ (9)
$\bm{q}\in\mathbb{R}^{d}$ is the query, $\bm{v}$ is the visited graph node, $\bm{w}$ is the neighbor of $\bm{v}$, and $\bm{e}=\bm{w}-\bm{v}$. $\tau$ is the threshold determined by the result list of graph. $\bm{q_{i}}$, $\bm{e_{i}}$ denote the $i$-th sub-vectors of $\bm{q}$ and $\bm{e}$, respectively ($1\leq i\leq L$). $\bm{u}^{i}_{j}$ denotes the $j$-th element of the $i$-th sub-vector of $\bm{u}$. $\bm{u^{i}_{e[i]}}$ denotes the reference vector of $\bm{e_{i}}$ among all $\{\bm{u^{i}_{j}}\}$’s ($1\leq j\leq m$). In our experiments, $S$ was set to $S_{\text{sym}}(256,L)$.
During the traversal of the similarity graph, we check the exact distance from graph node $w$ to $q$ only when Ineq. (9) is satisfied; otherwise, we skip the computation of $w$ for efficiency. A complete graph-based algorithm equipped with the KS2 test can be found in Alg. 6. By Lemma 4.3 and the same analysis in Lu et al. (2024), we can easily obtain the following result.
@@
Comparison with PEOs. Since PEOs also uses a Gaussian distribution to generate projection vectors in subspaces like CEOs, the estimation in Ineq. (9) is more accurate than that of the PEOs test, as discussed earlier. In addition, the proposed test has two advantages: (1) Ineq. (9) is much simpler than the testing inequality in the PEOs test, resulting in higher evaluation efficiency; (2) Ineq. (9) requires fewer constants to be stored, leading to a smaller index size compared to that of PEOs.
Complexity analysis. For the time complexity, for every edge $\bm{e}$, the computation of the LHS of Ineq. (9) requires $L$ lookups in the table and L1L-1 additions, while the computation of the RHS of Ineq. (9) requires two additions and one multiplication. By using SIMD, we can perform the KS2 test for $16$ edges simultaneously. For the space complexity, for each edge, we need to store $L$ bytes to recover $\bm{{q_{i}}^{\top}u^{i}_{e[i]}}$, along with two scalars, that is, $A_{S}(\bm{v})\|\bm{w}\|^{2}/(2\|\bm{e}\|)$ and $A_{S}(\bm{v})/\|\bm{e}\|$, which are quantized using scalar quantization to enable fast computation of the RHS of Ineq. (9).
@@ 7 Experiments
All experiments were conducted on a PC equipped with an Intel(R) Xeon(R) Gold 6258R CPU @ 2.70GHz. KS1 and KS2 were implemented in C++. The ANNS experiments used 64 threads for indexing and a single CPU for searching. We evaluated our methods on six high-dimensional real-world datasets: Word, GloVe1M, GloVe2M, Tiny, GIST, and SIFT. Detailed statistics for these datasets are provided in Appendix D.1. More experimental results can be found in Appendix D.

### v3

@@ 4 Analysis of Probability Guarantees
(2) Let $\bm{q}$, $\bm{v_{1}}$ and $\bm{v_{2}}$ be three normalized vectors on $\mathbb{S}^{d-1}$ such that $\left\langle\bm{q},\bm{v_{1}}\right\rangle>\left\langle\bm{q},\bm{v_{2}}\right\rangle$. The probability $\mathbb{P}[K^{1}_{S}(\bm{q},\bm{v_{1}})>K^{1}_{S}(\bm{q},\bm{v_{2}})|A_{S}(\bm{q})=\cos\psi]$ increases as $\psi$ decreases in $(0,\pi)$. In particular, when $\psi\in(0,\pi/2)$, $\mathbb{P}[K^{1}_{S}(\bm{q},\bm{v_{1}})>K^{1}_{S}(\bm{q},\bm{v_{2}})|A_{S}(\bm{q})=\cos\psi]>0.5$, that is, $K^{1}_{S}$ satisfies the probability guarantee in Problem 1.1.
@@ 5 Implementation and Complexity Analysis
We discuss how to configure $S$, and then analyze the complexity of $K^{1}_{S}$ and $K^{2}_{S}$. Based on the discussion in Sec. 4, we observe that small reference angles are preferred. Thus, given $m$, our goal is to construct a set $S$ of $m$ points on $\mathbb{S}^{d-1}$, denoted by $S_{m}$, such that $A_{S_{m}}(\cdot)$ is maximized. Due to the effect of the random rotation matrix $H$, the optimal configurations denoted by $\bar{S}_{m}$ and ${S}_{m}^{*}$, can be obtained either in the sense of expectation or in the sense of the worst case, respectively:
@@
where $U(\mathbb{S}^{d-1})$ denotes the uniform distribution on the sphere. By the definitions of $\bar{S}_{m}$ and ${S}_{m}^{\ast}$, they correspond to the configurations that achieve the largest expected value and the largest worst-case value of $A_{S}(\bm{v})$, respectively. On the other hand, finding the exact solutions for $\bar{S}_{m}$ and ${S}_{m}^{\ast}$ is closely related to the best covering problem, which is highly challenging and remains open in the general case. To the best of the authors’ knowledge, the optimal configuration ${S}_{m}^{\ast}$ is only known when $m\leq d+3$ (Borodachov et al., 2019). In light of this, we provide two configurations of $S$: one relies on random antipodal projections (Alg. 1), and the other is built using multiple cross-polytopes (Alg. 2). Each has its own advantages. Alg. 1 enables the estimation of reference angles, while Alg. 2 can empirically produce slightly smaller reference angles and is more efficient for projection computation.
@@
2   Generate $m/2$ points i.i.d. on $\mathbb{S}^{d^{\prime}-1}$, along with their antipodal counterparts  Scale the norm of all $m$ points in this iteration to $1/\sqrt{L}$, and collect the vectors after scaling
@@
Generate $N$ points randomly and independently on $\mathbb{S}^{d^{\prime}-1}$, where $N$ is a sufficiently large number
@@ 5      Repeat the above iteration and select $b/2$ antipodal pairs from the rotated cross-polytope
6  For the generated $S\in\mathbb{S}^{d^{\prime}-1}$, compute $\tilde{J}(S,N)$ and keep the configuration with the largest value, denoted by $S_{\text{max}}$
@@
Before proceeding into the details of algorithms, we introduce a quantity $J(S)$ as follows.
$J(S)=\mathbb{E}_{\bm{v}\in U(\mathbb{S}^{d-1})}[A_{S}(\bm{v})].$ (7)
By definition, $J(S)$ denotes the expected value of $A_{S}(\bm{v})$ w.r.t. $S$ (which equals the cosine of the reference angle when $\bm{v}\in\mathbb{S}^{d-1}$). This quantity is consistent with our theory, as a random rotation is applied to $\bm{v}$ or $\bm{q}$ in Eq. (2) and Eq. (3). Based on the previous discussion, for a fixed $m$, $J(S)$ is maximized when $S=\bar{S}_{m}$, which is hard to compute. Thus, we take $N$ to be sufficiently large and let $v_{1,N},\ldots,v_{N,N}$ be vectors drawn independently and uniformly from $U(\mathbb{S}^{d-1})$. Let $\tilde{J}(S,N)=[\sum_{i=1}^{N}A_{S}(\bm{v_{i,N}})]/N$. By the law of large numbers, we can approximate $J(S)$ by $\tilde{J}(S,N)$ when $N$ is sufficiently large.
@@
(1) (Utilization of antipodal pairs and cross-polytopes) We use the antipodal pair or the cross-polytope as our building block for the following three reasons. (i) Since all the projection vectors are antipodal pairs, the evaluation time of projection vectors can be halved. (ii) Both two structures can ensure that the assumption $A_{S}(\bm{v})>0$ holds, such that the condition in Lemma 4.3 is always satisfied. (iii) The result in Borodachov (2022) shows that, for $m=2d$, under mild conditions, the $2d$ vertices of a cross-polytope can be proven to have the smallest covering radius, that is, the smallest reference angle in the worst case. Although the results in the case $m>2d$ are unknown, we can rotate the fixed cross-polytope in random directions to generate multiple cross-polytopes until we obtain $m$ vectors, which explains the steps from 3 to 7 in Alg. 2.
@@
(3) (Accuracy boosting via multiple levels) Increasing $m$ can lead to a smaller reference angle. The analysis in Lu et al. (2024) shows that, for certain angle-thresholding problems requiring high accuracy, an exponential increase in $m$, rather than a linear one, can be effective. Therefore, similar to Lu et al. (2024), we use a product-quantization-like technique (Jégou et al., 2011) to partition the original space into $L$ subspaces (levels), which is adopted in Algs. 1 and 2. By concatenating equal-length sub-vectors from these $L$ subspaces, we can virtually generate $m^{L}$ normalized projected vectors. As will be shown in Lemma B.1, the introduction of $L$ can significantly decrease the reference angle, and $L$ can also control the trade-off between query accuracy and efficiency.
(4) (Fast projection computation via multi-cross-polytopes) In Eq. (2) and Eq. (3), we need to compute $HS\bm{v}$ in the indexing phase. If the projection time is a concern in practice, we can use the Fast Johnson–Lindenstrauss transformation to accelerate this process. Specifically, we use Alg. 2 with $R=1$. When $L=1$, the cost of computing $HS\bm{v}$ is $O(\max(d,m)\log d)$. When $L>1$, the cost of computing $HS\bm{v}$ can be reduced to $O(d\log d+mL\log(d/L))$.
@@
In the indexing phase, for a fixed dataset $\mathcal{D}$ of size $n$, the complexities of $K^{1}_{S}$ and $K^{2}_{S}$ are $O(nmd)$ and $O(nd\log d+nmd)$, respectively. In the query phase, $K^{1}_{S}$ requires $O(md)$- time for random projection, and for each $v\in\mathcal{D}$, it spends $O(1)$-time for computing $K^{1}_{S}(q,v)$, while $K^{2}_{S}$ spends $O(d\log d+md)$-time for random projection and random rotation, and for each $v\in\mathcal{D}$, it spends $O(L)$-time for computing $K^{2}_{S}(q,v)$, where $L\ll d$. Particularly, if $S=S_{\text{pol}}(m,L)$ with $R=1$, the complexities of $K^{1}_{S}$ and $K^{2}_{S}$ in the indexing phase can be reduced to $O(\max(d,m)n\log d)$ and $O(nd\log d+nmL\log(d/L))$, respectively, where $m$ is greater than $d/L$ by user-specification.
@@ 6.1 Improvement on CEOs-Based Techniques
As for $K^{1}_{S}$, we can use it to improve CEOs, which is used for MIPS and further applied to accelerate LSH-based ANNS (Andoni et al., 2015) and DBSCAN (Xu and Pham, 2024). Since CEOs was originally designed for inner products, we generalize $K^{1}_{S}$ to $K^{1^{\prime}}_{S}$ as follows to align with CEOs:
@@
It is easy to see that, with two minor modifications, that is, replacing $\bm{v}\in\mathbb{S}^{d-1}$ with $\bm{v}\in\mathbb{R}^{d}$, and replacing $x$ with $\|\bm{v}\|x$ in Eq. (4), Lemma 4.2 still holds. Therefore, $K^{1^{\prime}}_{S}$ can be regarded as a reasonable kernel function for inner products. Then, we can apply $K^{1^{\prime}}_{S}$ to the algorithm in Pham (2021); Pham and Liu (2022); Xu and Pham (2024). We only need to make the following modification. In these algorithms, the random Gaussian matrix, which denotes the set of projection vectors, can be replaced by $S_{\text{sym}}$ or $S_{\text{pol}}$, with the other parts unchanged. This substitution does not change the complexity of the original algorithms. To distinguish this projection technique based on $K^{1^{\prime}}_{S}$ from CEOs, we refer to it as KS1 (see Alg. 5 for the projection structure of KS1). In the experiments, we will demonstrate that KS1 yields a slight improvement in recall rates over CEOs, owing to a smaller reference angle.
@@ 6.2 A New Probabilistic Test in Similarity Graph
Lu et al. (2024) proposed a $(\delta,1-\epsilon)$-routing test called PEOs test. Based on $K^{2}_{S}$, we propose a new routing test for $\ell_{2}$ distance, called the KS2 test, as follows (see Sec. C.1 and Fig. 3 in Appendix for more details).
$\sum^{L}_{i=1}{\bm{{q_{i}}^{\top}u^{i}_{e[i]}}}\geq A_{S}(\bm{e})\cdot\frac{\|\bm{w}\|^{2}/2-\tau-\bm{v}^{\top}\bm{q}}{\|\bm{e}\|}.$ (9)
$\bm{q}\in\mathbb{R}^{d}$ is the query, $\bm{v}$ is the visited graph node, $\bm{w}$ is the neighbor of $\bm{v}$, and $\bm{e}=\bm{w}-\bm{v}$. $\tau$ is the threshold determined by the result list of graph search. $\bm{q_{i}}$, $\bm{e_{i}}$ denote the $i$-th sub-vectors of $\bm{q}$ and $\bm{e}$, respectively ($1\leq i\leq L$). $\bm{u}^{i}_{j}$ denotes the $j$-th sub-vector of the $i$-th subspace of $\bm{u}$. $\bm{u^{i}_{e[i]}}$ denotes the reference vector of $\bm{e_{i}}$ among all $\{\bm{u^{i}_{j}}\}$’s ($1\leq j\leq m$). In our experiments, $S$ was set to $S_{\text{sym}}(256,L)$.
During the traversal of the similarity graph, we check the exact distance from graph node $w$ to $q$ only when Ineq. (9) is satisfied; otherwise, we skip the computation of $w$ for efficiency. A complete graph-based algorithm equipped with the KS2 test can be found in Alg. 7. By Lemma 4.3 and the same analysis in Lu et al. (2024), we can easily obtain the following result.
@@
Comparison with PEOs. Since PEOs also uses a Gaussian distribution to generate projection vectors in subspaces, as CEOs does, and does not make use of the reference-angle information, the estimation in Ineq. (9) is more accurate than that of the PEOs test. In addition, the proposed test has two advantages: (1) Ineq. (9) is much simpler than the testing inequality in the PEOs test, resulting in higher evaluation efficiency; (2) Ineq. (9) requires fewer constants to be stored, leading to a smaller index size compared to that of PEOs.
Complexity analysis. For the time complexity, for every edge $\bm{e}$, the computation of the LHS of Ineq. (9) requires $L$ lookups in the table and L1L-1 additions, while the computation of the RHS of Ineq. (9) requires two additions and one multiplication. By using SIMD, we can perform the KS2 test for $16$ edges simultaneously. For the space complexity, for each edge, we need to store $L$ bytes to recover $\bm{{q_{i}}^{\top}u^{i}_{e[i]}}$, along with two scalars, that is, $A_{S}(\bm{e})\|\bm{w}\|^{2}/(2\|\bm{e}\|)$ and $A_{S}(\bm{e})/\|\bm{e}\|$, which are quantized using scalar quantization to enable fast computation of the RHS of Ineq. (9).
@@ 7 Experiments
All experiments were conducted on a PC equipped with an Intel(R) Xeon(R) Gold 6258R CPU @ 2.70GHz. KS1 and KS2 were implemented in C++. The ANNS experiments used 64 threads for indexing and a single CPU thread for searching. We evaluated our methods on six high-dimensional real-world datasets: Word, GloVe1M, GloVe2M, Tiny, GIST, and SIFT. Detailed statistics for these datasets are provided in Appendix D.1. More experimental results can be found in Appendix D.